OpenStudy (anonymous):

What is the parametric form of the linear equation 3x - 2y = 16

jhonyy9 (jhonyy9):

y=3x/2 -8

OpenStudy (anonymous):

This is the slope-intercept form, I need the parametric form.

sam (.sam.):

First you differentiate 3x - 2y = 16 \[3-2\frac{dy}{dx}=0\] \[3/2=\frac{dy}{dx}\]

sam (.sam.):

then you set \[\Huge \frac{\frac{dy}{dt}}{\frac{dx}{dt}}=\frac{3}{2}\] \[\Huge \frac{\frac{dy}{dt}}{\frac{dx}{dt}}=\frac{(\frac{3}{2})}{1}\] \[\frac{dy}{dt}=\frac{3}{2} ~\text{and}~ \frac{dx}{dt}=1\] Then integrate both dy/dt and dx/dt x=t+c , and y=3t/2 +d

OpenStudy (anonymous):

ty!

OpenStudy (rogue):

Differentiation and integration? No need to make it so complicated... 3x - 2y = 16 3x = 16 + 2y t = 3x, t = 16 + 2y x = t/3, y = .5t - 8